Radiation is contained in a volume V and has pressure P and total energy E .
Justify the correct value of PV
Answer: E/3
1
Expert's answer
2019-04-17T11:57:20-0400

Ideal gas equation can be written as

PV=nRT

nRT=NkT

Thus,

PV=NkT

Energy per atom we put on to the system

E=(3/2)kT

Total energy

E=N(3/2)kT

NkT=2/3E

Thus

PV=(2/3)E